# Foundation Models in Electrocardiogram: A Review

### v1

@@ 4 Discussion
Model weights are not shared. Out of 43 studies on ECG-FM, only 3 papers [10, 122, 67] have shared their saved weight files of the pre-trained model. These papers were able to do so because they utilized public datasets. Notably, the LLM community has a strong tradition of supporting and sharing their pre-trained FMs, as evidenced by the LLaMA-series, Baichuan-series, and Qwen-series models. However, the ECG research community lacks a clear guide on how to share pre-trained weights due to data privacy and ethics concerns, which hampers the development of ECG-FMs. Additionally, data isolation remains a significant challenge, with hospitals being hesitant to share their limited ECG data and collaborate to build a comprehensive ECG database. As a result, there is a lack of diversity in available ECG data for researchers and cardiologists. Furthermore, while high-tech companies typically develop and share state-of-the-art FMs, hospitals and medical institutions struggle to make full use of their real-time ECG data due to limited computing resources. Therefore, it is crucial to encourage collaboration among these institutions to drive the development of ECG-FMs and fully leverage valuable ECG data.

### v2

@@ 5 Discussion
Model weights are not shared. Out of 46 studies on ECG-FM, only 3 papers [10, 130, 72] have shared their saved weight files of the pre-trained model. These papers were able to do so because they utilized public datasets. Notably, the LLM community has a strong tradition of supporting and sharing their pre-trained FMs, as evidenced by the LLaMA-series, Baichuan-series, and Qwen-series models. However, the ECG research community lacks a clear guide on how to share pre-trained weights due to data privacy and ethics concerns, which hampers the development of ECG-FMs. Additionally, data isolation remains a significant challenge, with hospitals being hesitant to share their limited ECG data and collaborate to build a comprehensive ECG database. As a result, there is a lack of diversity in available ECG data for researchers and cardiologists. Furthermore, while high-tech companies typically develop and share state-of-the-art FMs, hospitals and medical institutions struggle to make full use of their real-time ECG data due to limited computing resources. Therefore, it is crucial to encourage collaboration among these institutions to drive the development of ECG-FMs and fully leverage valuable ECG data.